A contractor has purchased a wheel loader for $115,000 and plans to use it 2,000 hours per year. The cost of one set of tires is $25,000. At this usage rate, the contractor anticipates disposing of the loader after using it for 10 years and realizing a salvage value of $35,000. The flywheel horsepower rating of the loader’s diesel engine is 105 horsepower. The interest rate is 10%. The loader operator will earn $34.00 per hour including fringe benefits, and diesel fuel costs $1.20 per gallon.

How much is the contractor’s hourly depreciation portion of ownership cost for the loader if using the Average Annual Investment Method?
Answer #1

Ans:

Initial price = 115,000
Cost of tires =25,000

Purchase price less tires = 90,000

Uniform series required to replace 90,000 present value

= 90,000 X (10% (1+10%)^10) / (1+10%)-1 = 90,000 X 0.162745 = 14,647.09

sinking factor to use with salvage value

= 50,000 X (10%/(1+10%)-1) = 50,000 X 0.062745 = 2196.089

Portion of ownership cost = (14,647 - 2,196) / 2000 hour per year = 6.23